# Ekpyrosis in Quantum Gravitational Anisotropic Bouncing Models

### v1

@@ II.2 Effective dynamics of Bianchi-I model
Moreover, it has been shown in [26], that strong curvature singularities are generically resolved and the spacetime is geodesically complete in the effective Bianchi-I model. The bounce in the Bianchi-I model is often accompanied by Kasner transitions in the geometry of the spacetime, as shown by extensive numerical simulations [36]. Note that, unlike classical dynamics, $\sigma^{2}$ no longer has a simple dependence on the mean scale factor $a$, and the modified generalized Friedmann and Raychaudhuri equations are as yet unknown for the effective dynamics of the Bianchi-I model. In [59], an interesting parabolic relation between the energy density and the shear scalar was numerically shown to hold at the quantum bounce.
@@ II.3 Classical dynamics of Bianchi-IX model
[…]
@@ II.4 Effective Dynamics of Bianchi-IX model
Note that this is the exact same bound as found in the Bianchi-I effective spacetime in equation (20), as this expression may be derived from the directional Hubble rates using equation (18).
@@
In the following sections, we perform the numerical simulations of effective dynamics for Bianchi-I and Bianchi-IX spacetimes. The initial conditions are chosen in the contracting branch far from the bounce where the universe is macroscopic and the effective spacetime is well-approximated by classical dynamics. Apart from the matter content, we have six phase space variables in both Bianchi-I and Bianchi-IX models, and one constraint; namely the Hamiltonian constraint. For simulations, we first choose the initial values of $c_{1},c_{2},p_{1},p_{2},p_{3},\phi$ and $\rho$. These are chosen randomly from a uniform distribution over the given ranges: $c_{1},c_{2}\in[-0.25,0]$, $p_{1},p_{2},p_{3}\in[10000,30000]$, $\phi\in[0,0.4]$ and $\rho\in[0,10^{-4}]$. We then solve the Hamiltonian constraint to obtain the initial value of $c_{3}$. The initial value of $p_{\phi}$ can be obtained from the initial values of $\phi$ and $\rho$. The effective equations of motion for the triad and connection variables described earlier are then used to obtain the time evolution. For both Bianchi-I and Bianchi-IX models, $500$ simulations each are performed for both ekpyrotic and ekpyrotic-like potentials as well as for the massless scalar field for comparison. When compared to the massless case, the values of $c_{1}$, $c_{2}$, all triads and $\phi$ and $\rho$ were taken as same.
@@ III.1 Ekpyrotic potential
Figure 1: An example of the non-singular evolution of directional scale factors, equation of state, energy density and anisotropic shear in the bounce regime for effective Bianchi-I spacetime coupled to the ekpyrotic scalar field with $U_{0}=0.25$ in (41). It is typical to get multiple bounces before the universe eventually enters a macroscopic expanding regime. Unlike the classical theory, the scale factors have finite nonzero values throughout the evolution and the energy density and anisotropic shear do not diverge. These features are seen in all simulations discussed in this manuscript.
@@ III.2 Ekpyrotic-like potential
Fig. 6 shows the comparison of the energy density and shear scalar at the last bounce with their respective values in the case of a massless scalar field for $U_{0}=0.25$. We find that as in the case of the ekpyrotic potential considered above, the ekpyrotic-like potential also leads to isotropization at the bounce. In particular, the energy density at the bounce is increased compared to the massless case in all but a few simulations. However, this is not necessarily accompanied by a decrease in the shear scalar, as seen by the color coding in the energy density plot in Fig. 6 where only the red points experienced $\sigma_{\mathrm{b}}^{2}$ decrease. The second panel on the comparison of the shear scalar confirms the same, although $61.00\%$ of the simulations did have $\sigma_{\mathrm{b}}^{2}$ decrease compared to a massless scalar. Yet, the relative strength of the anisotropies, represented by the ratio $R_{\sigma}$ at the bounce, is decreased by the ekpyrotic-like potential in all but a few cases, as seen in the third panel of Fig. 6. Moreover, similar to the case of effective Bianchi-I spacetime with ekpyrotic potential, there is no mixing of the red and blue dots in the energy density plots indicating a clear boundary between the two populations of simulations.
@@ IV.2 Ekpyrotic-like potential
Table 1: Comparison of changes in $\rho$, $\sigma^{2}$ for matter fields with different potentials (the numbers represent the magnitude of $\rho$ and $\sigma^{2}$ at the bounce as a fraction of their respective maximum possible value in LQC). $R_{\sigma}$ is a measure of the relative strength of anisotropic shear with respect to energy density. The values of these parameters averaged over all 500 simulations for a given spacetime and potential configuration are $\rho_{\mathrm{av}},\sigma^{2}_{\mathrm{av}}$ and $R_{\sigma_{\mathrm{av}}}$. Note that ekpyrotic and ekpyrotic-like potential with the same value of $U_{o}$ are not directly comparable as their corresponding potential strengths are quite different for the same values of $U_{0}$.
@@ V Summary
Bouncing cosmologies where quantum gravitational effects lead to singularity resolution have been explored as potential scenarios to explain the conditions in the very early universe and are often discussed as potential alternatives to inflation. Earlier phenomenological studies in LQC strongly indicate that in the presence of ordinary matter, the Bianchi-I and Bianchi-IX effective spacetimes remain highly anisotropic after the bounce. Moreover, for such matter the bounce is more likely to increase anisotropies as anisotropies tend to dominate over matter in the bounce regime. This is primarily due to the fact that anisotropies grow at a much faster rate than ordinary matter when a contracting universe approaches the bounce. Thus, additional input is needed if isotropization is to be achieved. In our analysis, we considered whether an ekpyrotic scalar field having an ultra-stiff equation of state, may not only help curb the growth of anisotropies but also enhance the isotropization of the resulting universe. Specifically, using the anisotropic bouncing cosmologies of LQC, we studied the problem of isotropization at the bounce with the help of two different ekpyrotic potentials, first in the effective Bianchi-I spacetime as the simplest anisotropic scenario, and then in effective Bianchi-IX spacetimes as the most generic anisotropic scenario including spatial curvature. The intuitive reasons for expecting the ekpyrotic field to produce isotropization are easy to understand. While anisotropies in the bounce regime have an effective equation of state equivalent to that of stiff matter, the ekpyrotic field behaves as an isotropic fluid having an equation of state larger than unity. Thus, the ekpyrotic field is expected to grow faster than anisotropies as the universe contracts towards a bounce and produces isotropization [43]. Recently, a detailed analysis of the ekpyrotic scenario in LQC was carried out in [50], where it was found that although the presence of an ekpyrotic field-dominated regime with a phase of equation of state larger than unity in the vicinity of the bounce was a robust feature of the background dynamics, its duration and strength was rather small. This begs the question, will the ekpyrotic field be effective in causing isotropization of the bounce? To explore this, we consider the anisotropic Bianchi-I and Bianchi-IX models in LQC in the effective spacetime description as our setting to put the isotropizing ability of the ekpyrotic field to the test. Specifically, we study whether the ekpyrotic field can dominate the bounce regime and lead to a reduction in the strength of the anisotropic shear relative to the energy density of the universe (which is taken to be isotropic in this manuscript).
@@
Through extensive numerical simulations, we studied the effect of ekpyrotic and ekpyrotic-like potentials and found that they have a strong isotropization effect on the universe. We found that compared to the massless scalar field as the matter content of the universe, the inclusion of these potentials generically decreased $R_{\sigma}$, our measure of anisotropy which indicates the relative strength of the anisotropic shear compared to the isotropic energy density, in over $90\%$ of cases; regardless of which effective spacetime or potential was considered. This indicates that these potentials do in fact have a strong isotropization effect. Further, as found in above mentioned previous studies, we observe that the phase of ekpyrosis near any bounce is always found to be short-lived. However, these short-lived phases of ekpyrosis coupled with multiple bounces are sufficient to cause the isotropization observed in this manuscript. Importantly, our simulations indicate that this reduction in $R_{\sigma}$ is primarily achieved by an enhancement of the isotropic energy density at the bounce. Thus, while the energy density at the bounce is increased compared to massless scalar field in almost all cases, the effect on the anisotropic shear is not as decisive.
@@
A comparison of the effects of the two ekpyrotic potentials in Bianchi-I and Bianchi-IX effective spacetimes is given in Table 1. For generic initial conditions, the simulations often resulted in a succession of bounces before the emergence of a semi-classical universe. Since we are interested in isotropization at the bounce as a mechanism for isotropization of the large-scale universe, we attempted to pick time intervals of simulations so that a semi-classical expanding universe is finally obtained, and the results reported in Table 1 are for the last bounce. However, the extremely complex behavior of the dynamical equations of the effective spacetime, particularly for Bianchi-IX universe, led to difficulties in choosing such a time scale for the simulations. We also studied the effect of increasing the strength of the potentials, indicated by the potential parameter $U_{0}$. As shown in the table, increasing the strength of the potential often led to higher isotropization, though not always. We also note that isotropization is achieved in more cases in the Bianchi-I spacetime compared to the Bianchi-IX spacetime. This is expected as the presence of spatial curvature in the Bianchi-IX spacetime complicates the interplay between anisotropic shear and energy density.

### v2

@@ II.2 Effective dynamics of Bianchi-I model
Moreover, it has been shown in [27], that strong curvature singularities are generically resolved and the spacetime is geodesically complete in the effective Bianchi-I model. The bounce in the Bianchi-I model is often accompanied by Kasner transitions in the geometry of the spacetime, as shown by extensive numerical simulations [36]. These numerical simulations show that even though $\rho_{\mathrm{max}}$ and $\sigma^{2}_{\mathrm{max}}$ are theoretical maxima and due to the joint contribution of energy density and anisotropic shear to the spacetime curvature, the singularity resolution via bounces can occur at values smaller than these upper bounds. Note that, unlike classical dynamics, $\sigma^{2}$ no longer has a simple dependence on the mean scale factor $a$, and the modified generalized Friedmann and Raychaudhuri equations are as yet unknown for the effective dynamics of the Bianchi-I model. In [63], an interesting parabolic relation between the energy density and the shear scalar was numerically shown to hold at the quantum bounce.
@@ II.3 Classical dynamics of Bianchi-IX model
$p_{1}=\mathrm{sgn}(a_{1})|a_{2}a_{3}|{l_{o}}^{2},\quad\text{$p_{2}$ and $p_{3}$ given by cyclic permutations},.$ (26)
@@ II.4 Effective Dynamics of Bianchi-IX model
Note that this is the exact same bound as found in the Bianchi-I effective spacetime in equation (20), as this expression may be derived from the directional Hubble rates using equation (18). As in the case of Bianchi-I model in LQC, the bounce in Bianchi-IX spacetime can occur at energy density and anisotropic shear at values below their maximum allowed values.
@@
In the following sections, we perform the numerical simulations of effective dynamics for Bianchi-I and Bianchi-IX spacetimes. The initial conditions are chosen in the contracting branch far from the bounce where the universe is macroscopic and the effective spacetime is well-approximated by classical dynamics. Apart from the matter content, we have six phase space variables in both Bianchi-I and Bianchi-IX models, and one constraint; namely the Hamiltonian constraint. For simulations, we first choose the initial values of $c_{1},c_{2},p_{1},p_{2},p_{3},\phi$ and $\rho$. The initial conditions chosen for all the simulations in the manuscript correspond to a classical macroscopic universe with energy density positive. The value of connection component $c_{3}$ is determined using the vanishing of Hamiltonian constraint. The values of $c_{1}$ and $c_{2}$ are chosen in range $c_{1},c_{2}\in[-0.25,0]$. The values of triads are chosen in range $p_{1},p_{2},p_{3}\in[10000,30000]$, and the value of initial energy density is chosen in the range $\rho\in[0,10^{-4}]$. The value of scalar field was chosen in the range $\phi\in[0,0.4]$. The initial value of $p_{\phi}$ can be obtained from the initial values of $\phi$ and $\rho$. [^4] The actual allowed range of $\phi$ varies with the potential parameters. As an example, for the ekpyrotic potential (41) with $U_{o}=0.25$, $\sigma_{1}=0.3\sqrt{8\pi}$ and $\sigma_{2}=0.09\sqrt{8\pi}$, $\phi$ can not be more than 0.0007 for $\rho=10^{-4}$. Using the relationship between $p_{\phi}$ and $\dot{\phi}$, the range of $\dot{\phi}$ falls roughly between $\dot{\phi}\in[0,1.5\times 10^{-2}]$ for the ekpyrotic potential and $\dot{\phi}\in[0,1]$ for the ekpyrotic-like potential. The effective equations of motion for the triad and connection variables described earlier are then used to obtain the time evolution. For both Bianchi-I and Bianchi-IX models, $500$ simulations each are performed for both ekpyrotic and ekpyrotic-like potentials as well as for the massless scalar field for comparison. When compared to the massless case, the values of $c_{1}$, $c_{2}$, all triads and $\phi$ and $\rho$ were taken as same. As discussed earlier in the Introduction, recall that we are considering a macroscopic contracting universe in which the ekpyrotic field is initially in the positive regime moving towards smaller $\phi$ values. While ekpyrosis can occur when the field enters the negative well, the initial conditions considered in our work do not assume existence of ekpyrosis.
Note that the purpose of this work is not to obtain viable ekpyrotic cosmology per se, rather it is to study the effect of the ekpyrotic potential on anisotropies starting from generic initial conditions corresponding to macroscopic contracting universe. Further, the initial conditions are not specifically tailored to obtain a definite phase of ekpyrosis (a phase with $w\gg 1$), instead the existence and duration of the ekpyrosis is left to be determined by the dynamics while setting random initial conditions for the scalar field from a chosen range of values corresponding to a macroscopic contracting universe. In other words, an extended phase of ekpyrosis is not guaranteed by the initial conditions chosen in our numerical simulations. We have chosen positive initial values for the scalar field in this study, wherein the permitted initial range of values of $\phi$ and $\dot{\phi}$ are determined by the fact that we set our initial conditions in the far past where the universe is large and contracting and the total energy density is very small compared to the Planck scale. Further, as noted above, in LQC there are also global bounds on energy density and shear scalar determined by the underlying quantum geometry. It is also important to note that at the bounce these theoretical maximum values are generally not saturated and singularity resolution can occur in anisotropic spacetime at values lower than $\rho_{\mathrm{max}}$ and $\sigma^{2}_{\mathrm{max}}$. This produces a constraint on the allowed initial conditions by the Hamiltonian constraint in loop quantum cosmology. As an example, it is not possible to consider the initial value of $\phi$ of the order unity and large $\dot{\phi}$ in our numerical simulations for the ekpyrotic potential. For concreteness, consider the case of the ekpyrotic potential (41) with $U_{o}=0.25$, $\sigma_{1}=0.3\sqrt{8\pi}$ and $\sigma_{2}=0.09\sqrt{8\pi}$ in Bianchi-I spacetime as an example. If one takes $\phi\approx 1$ as the initial condition, then $U\approx 0.1$ in Planck units. Even if the kinetic energy of the field is vanishing, this contribution from potential energy means that the universe already is in the Planck regime in terms of the value of energy density. Thus, such an initial condition is not compatible with an initial macroscopic classical universe with a small energy density. Since the maximum bound on energy density is 0.41, this means that the maximum allowed contribution from kinetic energy term is 0.31 in Planck units. Even if one assumes that energy density at the bounce reaches the maximum theoretical value, one finds that the magnitude of $\dot{\phi}$ is approximately 0.8. Therefore, one can not have initial conditions with $\phi$ of the order unity and $\dot{\phi}$ very large in Planck units in LQC. In contrast, in case of the ekpyrotic-like potential which is negative definite, it is indeed possible to consider large and positive initial values of $\phi$. However, the ekpyrotic-like potential already assumes very small values in the range of initial values considered in this manuscript, i.e. $\phi\in[0,0.4]$. The ekpyrotic-like potential keeps decreasing in magnitude gradually for higher values of $\phi$, such that going to values higher than the considered range is unlikely to yield radically different results than those obtained in this manuscript. Note that both the potentials have a single global minimum adjacant to $\phi=0$. While the ekpyrotic-like potential remains negative and approaches vanishing values on either side of the minimum, the ekpyrotic potential becomes positive on the right branch of the minima beyond $\phi=0$. Our initial conditions for the field are set on the right edge of the potential minimum with the field evolving towards the minimum as the universe contracts. While the initial conditions considered in this manuscript result in short-lived phases of ekpyrosis, we still find significant isotropization.
@@ III.1 Ekpyrotic potential
Figure 1: An example of the non-singular evolution of directional scale factors, equation of state, energy density and anisotropic shear in the bounce regime for effective Bianchi-I spacetime coupled to the ekpyrotic scalar field with $U_{0}=0.25$ in (41). If one considers time from negative to positive three scale factors are decreasing before $t=0$ and then two are increasing while one is decreasing after the bounce. The universe has a point like approach to singularity which after resolution due to quantum geometry effects yields a cigar like evolution as is generally the case for Kasner transitions in LQC [36]. It is typical to get multiple bounces before the universe eventually enters a macroscopic expanding regime. Unlike the classical theory, the scale factors have finite nonzero values throughout the evolution and the energy density and anisotropic shear do not diverge. These features are seen in all simulations discussed in this manuscript.
@@ III.2 Ekpyrotic-like potential
Fig. 6 shows the comparison of the energy density and shear scalar at the last bounce with their respective values in the case of a massless scalar field for $U_{0}=0.25$. We find that as in the case of the ekpyrotic potential considered above, the ekpyrotic-like potential also leads to isotropization at the bounce. Note that in comparison to the simulation for the ekpyrotic potential shown in Fig. 1, the bounce appears smoother. Though, there is no general reason for the ekpyrotic-like potential to give a smoother bounce, the contrast between Fig. 1 and Fig. 5 is due to the difference in initial conditions, especially of the potential and the resulting impact on the equation of state. In the ekpyrotic-like case shown here the potential is about 100 times larger in magnitude than the case shown in Fig. 1 for a similar value of initial energy density. Further, the the initial equation of state at $t=0$ is highly stiff ($w\approx 13.86$) compared to the ekpyrotic potential case in Fig. 1 where it is $w\approx 0.82$. We suspect that the initial equation of state being highly stiff causes the bounce to be more isotropic and smoother than the case corresponding to Fig. 1.
In the ekpyrotic-like case, the energy density at the bounce is increased compared to the massless case in all but a few simulations. However, this is not necessarily accompanied by a decrease in the shear scalar, as seen by the color coding in the energy density plot in Fig. 6 where only the red points experienced $\sigma_{\mathrm{b}}^{2}$ decrease. The second panel on the comparison of the shear scalar confirms the same, although $61.00\%$ of the simulations did have $\sigma_{\mathrm{b}}^{2}$ decrease compared to a massless scalar. Yet, the relative strength of the anisotropies, represented by the ratio $R_{\sigma}$ at the bounce, is decreased by the ekpyrotic-like potential in all but a few cases, as seen in the third panel of Fig. 6. Moreover, similar to the case of effective Bianchi-I spacetime with ekpyrotic potential, there is no mixing of the red and blue dots in the energy density plots indicating a clear boundary between the two populations of simulations.
@@ IV.2 Ekpyrotic-like potential
Table 1: The table summarizes results from various numerical simulations for Bianchi-I and Bianchi-IX spacetime in LQC. 500 simulations were performed for each case. Comparison of changes in $\rho$, $\sigma^{2}$ for matter fields with different potentials (the numbers represent the magnitude of $\rho$ and $\sigma^{2}$ at the bounce as a fraction of their respective maximum possible value in LQC). $R_{\sigma}$ is a measure of the relative strength of anisotropic shear with respect to energy density. The initial values of $c_{1}$ and $c_{2}$ are chosen in range $c_{1},c_{2}\in[-0.25,0]$, whereas the values of triads are chosen in range $p_{1},p_{2},p_{3}\in[10000,30000]$. The value of $c_{3}$ is determined using the Hamiltonian constraint. The value of initial energy density is chosen in the range $\rho\in[0,10^{-4}]$. For $U_{o}=0.25$, the range of $\phi$ lies in the range $\phi\in[0,0.0007]$ for the ekpyrotic potential (41), whereas $\phi\in[0,0.4]$ for the ekpyrotic-like potential (43). Also $\dot{\phi}\in[0,1.5\times 10^{-2}]$ for the ekpyrotic potential and $\dot{\phi}\in[0,1]$ for the ekpyrotic-like potential. These values change slightly for $U_{o}=1.25$. The values of $\rho$, $\sigma^{2}$ and $R_{\sigma}$ averaged over all 500 simulations for a given spacetime and potential configuration are denoted as $\rho_{\mathrm{av}},\sigma^{2}_{\mathrm{av}}$ and $R_{\sigma_{\mathrm{av}}}$. Note that ekpyrotic and ekpyrotic-like potential with the same value of $U_{o}$ are not directly comparable as their corresponding potential strengths are quite different for the same values of $U_{0}$.
$\phi\in[0,0.4]$ for the ekpyrotic-like potential (43).
@@ V Summary
Bouncing cosmologies where quantum gravitational effects lead to singularity resolution have been explored as potential scenarios to explain the conditions in the very early universe and are often discussed as potential alternatives to inflation. Earlier phenomenological studies in LQC strongly indicate that in the presence of ordinary matter, the Bianchi-I and Bianchi-IX effective spacetimes remain highly anisotropic after the bounce. Moreover, for such matter the bounce is more likely to increase anisotropies as anisotropies tend to dominate over matter in the bounce regime. This is primarily due to the fact that anisotropies grow at a much faster rate than ordinary matter when a contracting universe approaches the bounce. Thus, additional input is needed if isotropization is to be achieved. In our analysis, we considered whether an ekpyrotic scalar field having an ultra-stiff equation of state, may not only help curb the growth of anisotropies but also enhance the isotropization of the resulting universe. Specifically, using the anisotropic bouncing cosmologies of LQC, we studied the problem of isotropization at the bounce with the help of two different ekpyrotic potentials, first in the effective Bianchi-I spacetime as the simplest anisotropic scenario, and then in effective Bianchi-IX spacetimes as the most generic anisotropic scenario including spatial curvature. The intuitive reasons for expecting the ekpyrotic field to produce isotropization are easy to understand. While anisotropies in the bounce regime have an effective equation of state equivalent to that of stiff matter, the ekpyrotic field behaves as an isotropic fluid having an equation of state larger than unity. Thus, the ekpyrotic field is expected to grow faster than anisotropies as the universe contracts towards a bounce and produces isotropization [47]. Analysis of the singularity resolution coupled to ekpyrotic scenario in LQC has been carried out earlier, but the question whether ekpyrotic field is effective in causing isotropization of the bounce was not addressed. To explore this, we consider the anisotropic Bianchi-I and Bianchi-IX models in LQC in the effective spacetime description as our setting to put the isotropizing ability of the ekpyrotic field to the test. Specifically, we study whether the ekpyrotic field can dominate the bounce regime and lead to a reduction in the strength of the anisotropic shear relative to the energy density of the universe (which is taken to be isotropic in this manuscript). It is important to note that the initial conditions are not specifically geared towards achieving an extended phase of ekpyrosis. The randomized initial conditions are chosen to be agnostic towards the occurrence and duration of a phase of ekpyrosis, which is left to be determined entirely by the underlying cosmological dynamics and the interplay between energy density and anisotropic shear. The objective of this work is to study how likely the ekpyrotic field is to overwhelm the anisotropic shear, starting from random initial conditions given in the classical macroscopic phase of the universe which is contracting. Thus, the initial conditions are not deliberately made favorable towards the ekpyrotic field.
@@
Through extensive numerical simulations, we studied the effect of ekpyrotic and ekpyrotic-like potentials and found that they have a strong isotropization effect on the universe. We found that compared to the massless scalar field as the matter content of the universe, the inclusion of these potentials generically decreased $R_{\sigma}$, our measure of anisotropy which indicates the relative strength of the anisotropic shear compared to the isotropic energy density, in over $90\%$ of cases; regardless of which effective spacetime or potential was considered. It should be noted that in our numerical simulations, the initial conditions of the scalar field and its velocity were small which are not ideal for ekpyrosis to occur. This is the consequence of theoretical upper bounds on energy density and anisotropic shear in LQC. Even with such not favorable initial conditions we find that sufficient isotropization is achieved. We find that even short-lived phases of ekpyrosis coupled with multiple bounces are sufficient to cause the isotropization for the initial conditions considered in our work. This indicates that for a different quantization scheme allowing larger bounds on energy density and anisotropic shear, the ekpyrotic phase is expected to be stronger. Importantly, our simulations indicate that this reduction in $R_{\sigma}$ is primarily achieved by an enhancement of the isotropic energy density at the bounce. Thus, while the energy density at the bounce is increased compared to massless scalar field in almost all cases, the effect on the anisotropic shear is not as decisive.
@@
A comparison of the effects of the two ekpyrotic potentials in Bianchi-I and Bianchi-IX effective spacetimes is given in Table 1. The simulations often resulted in a succession of bounces before the emergence of a semi-classical universe. Since we are interested in isotropization at the bounce as a mechanism for isotropization of the large-scale universe, we attempted to pick time intervals of simulations so that a semi-classical expanding universe is finally obtained, and the results reported in Table 1 are for the last bounce. However, the extremely complex behavior of the dynamical equations of the effective spacetime, particularly for Bianchi-IX universe, led to difficulties in choosing such a time scale for the simulations. We also studied the effect of increasing the strength of the potentials, indicated by the potential parameter $U_{0}$. As shown in the table, increasing the strength of the potential often led to higher isotropization, though not always. We also note that isotropization is achieved in more cases in the Bianchi-I spacetime compared to the Bianchi-IX spacetime. This is expected as the presence of spatial curvature in the Bianchi-IX spacetime complicates the interplay between anisotropic shear and energy density.